The mean of the test scores of a class of n students is 68 and the mean of the
test scores of a class of m students is 94.The mean score, when both classes are combined, is 88.What is the value of m/n ?

Answer by stanbon(57328)

The mean of the test scores of a class of n students is 68
Total points for those n students is 68n
-----------------------
and the mean of the test scores of a class of m students is 94.
Total points for those m students is 94m
--------
The mean score, when both classes are combined, is 88.What is the value of m/n ?
---
Equation:
(68n+94m)/(n+m) = 88
---
68n + 94m = 88n + 88m
20n = 6m
m/n= 20/6 = 10/3
=======================
Cheers,
Stan H.